OpenStudy (anonymous):

PLEASE HELP is this a legitimate funcion? f(x)=11x+5 HOW DO I SOLVE IT????????!!?!?!?!?!

OpenStudy (anonymous):

you don't it's a function, not an equation.

jimthompson5910 (jim_thompson5910):

Sure it's a legitimate function because each input you plug in for x produces exactly one output for y or f(x) As for solving it, are you asking how to find the x-intercept?

jimthompson5910 (jim_thompson5910):

To find the inverse, you replace f(x) with y. Then you swap x and y. Then solve for y. \[\Large f(x)=11x+5\] \[\Large y=11x+5\] \[\Large x=11y+5\] \[\Large x-5=11y\] \[\Large 11y=x-5\] \[\Large y=\frac{x-5}{11}\] \[\Large f^{-1}(x)=\frac{x-5}{11}\]

jimthompson5910 (jim_thompson5910):

If you want to evaluate f(3), then replace x with 3 in f(x) to get... f(x) = 11x + 5 f(3) = 11(3) + 5 f(3) = 33 + 5 f(3) = 38

jimthompson5910 (jim_thompson5910):

Evaluating \(\Large f^{-1}(3)\) is done the same way (but using the inverse this time) \[\Large f^{-1}(x)=\frac{x-5}{11}\] \[\Large f^{-1}(3)=\frac{3-5}{11}\] \[\Large f^{-1}(3)=-\frac{2}{11}\]

jimthompson5910 (jim_thompson5910):

sure thing the last bit you need to do is plug in any number for x into f(x) then take that result and plug it into the inverse

jimthompson5910 (jim_thompson5910):

I'm glad things are clicking now